Solve using any method: 6x + 4y = 34 - 9.2 - 6 - 51
Answer #1

Given that,

6x + 4y = 34...................(1)

and, -9x - 6y = -51.............(2)

.

Now 6*(1) + 4*(2) gives,

36x + 24y - 36x - 24y = 204 - 204

i.e. 0 = 0

.

This implies that both the equations are same which implies that there are infinitely many solutions.

.

Ans the solution is of the form,

\LARGE \boldsymbol{(x, y)= (x, \frac{34-6x}{4})}